# Agent Security Bench (ASB): Formalizing and Benchmarking Attacks and Defenses in LLM-based Agents

### v1

@@ Abstract
Although LLM-based agents, powered by Large Language Models (LLMs), can use external tools and memory mechanisms to solve complex real-world tasks, they may also introduce critical security vulnerabilities. However, the existing literature does not comprehensively evaluate attacks and defenses against LLM-based agents. To address this, we introduce Agent Security Bench (ASB), a comprehensive framework designed to formalize, benchmark, and evaluate the attacks and defenses of LLM-based agents, including 10 scenarios (e.g., e-commerce, autonomous driving, finance), 10 agents targeting the scenarios, over 400 tools, 23 different types of attack/defense methods, and 8 evaluation metrics. Based on ASB, we benchmark 10 prompt injection attacks, a memory poisoning attack, a novel Plan-of-Thought backdoor attack, a mixed attack, and 10 corresponding defenses across 13 LLM backbones with nearly 90,000 testing cases in total. Our benchmark results reveal critical vulnerabilities in different stages of agent operation, including system prompt, user prompt handling, tool usage, and memory retrieval, with the highest average attack success rate of 84.30%, but limited effectiveness shown in current defenses, unveiling important works to be done in terms of agent security for the community. Our code can be found at https://github.com/agiresearch/ASB.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced in their capabilities, enabling them to perform tasks such as content generation, question answering, tool calling, coding and many others (Kojima et al., 2024; Huang et al., 2022). This has paved the way for developing AI agents that combine LLMs with tools and memory mechanisms capable of interacting with broader environments (Ge et al., 2023a). These LLM-based agents have the potential to be deployed in various roles, such as safety-critical domains like financial services (Yu et al., 2023b), medical care (Abbasian et al., 2024; Yang et al., 2024b), and autonomous driving (Mao et al., 2024). As shown in Fig. 1, an LLM-based agent based on ReAct framework (Yao et al., 2022) usually operates through several key steps when solving a task: ① Defining roles and behaviors via a system prompt. ② Receiving user instructions and task details. ③ Retrieving relevant information from a memory database. ④ Planning based on the retrieved information and prior context. ⑤ Executing actions using external tools.
@@
Although recent research on LLM agents and advanced frameworks has made significant progress, the primary emphasis has been on their effectiveness and generalization (Qin et al., 2024; Mei et al., 2024; Ge et al., 2023b), with their trustworthiness remaining largely under-investigated (Hua et al., 2024). Specifically, while each of these steps mentioned above enables the agent to perform highly complex tasks, they also provide attackers with multiple points of access to compromise the agent system. Each stage is vulnerable to different types of adversarial attacks. Although several benchmarks have been proposed to evaluate the security of LLM agents, such as InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024), they are often limited by their scope, assessing either a single type of attack, i.e., indirect prompt injection, or operating in only a few scenarios, such as financial harm and data security. To address these limitations, we introduce Agent Security Bench (ASB), a comprehensive benchmark that formalizes and evaluates a wide range of adversarial attacks and defenses on LLM-based agents in ten different scenarios.
@@
Primarily, ASB covers various attack and defense types targeting each operational step of LLM-based agents, including system prompt, user prompt handling, tool usage, and memory retrieval. It evaluates Direct Prompt Injections (DPI), Observation Prompt Injections (OPI), Memory Poisoning, Plan-of-Thought (PoT) Backdoor Attacks, Mixed Attacks, and their defenses, offering the first holistic assessment of LLM agents’ security. In detail, a straightforward way to compromise an agent is through DPI, where attackers directly manipulate the user prompt to guide the agent toward malicious actions. Additionally, the agent’s reliance on external tools introduces further risks, particularly as attackers can embed harmful instructions into tool responses, referred to as OPI. Moreover, the planning phase of LLM agents faces security risks, as long-term memory modules like RAG databases (Lewis et al., 2020) can be compromised by memory poisoning attacks, where adversaries inject malicious task plans or instructions to mislead the agent in future tasks. In addition, since the system prompt is typically hidden from the user, it becomes a tempting target for Plan-of-Thought (PoT) Backdoor Attacks, where attackers embed hidden instructions into the system prompt to trigger unintended actions under specific conditions. Finally, attackers can also combine them to create mixed attacks that target multiple vulnerabilities across different stages of the agent’s operation.
@@
Our key contributions are summarized as follows: ① We design and develop Agent Security Bench (ASB), the first comprehensive benchmark including 10 scenarios (e.g., e-commerce, autonomous driving, finance), 10 agents targeting the scenarios, over 400 tools and tasks for evaluating the security of LLM-based agents against numerous attacks and defense strategies. ② We propose a novel PoT Backdoor Attack, which embeds hidden instructions into the system prompt, exploiting the agent’s planning process to achieve high attack success rates. ③ We formalize and categorize various adversarial threats targeting key components of LLM agents, including DPI, OPI, Memory Poisoning Attacks, PoT Backdoor Attacks, and Mixed Attacks, covering vulnerabilities in system prompt definition, user prompt handling, memory retrieval, and tool usage. ④ We benchmark 23 different types of attacks and defenses on ASB across 13 LLM backbones, demonstrating that LLM-based agents are vulnerable to the attacks, with the highest average attack success rates exceeding 84.30%. In contrast, existing defenses are often ineffective. Our work highlights the need for stronger defenses to protect LLM agents from sophisticated adversarial techniques.
Figure 1: Overview of the LLM Agent Attacking Framework, including Direct Prompt Injections (DPI), Observation Prompt Injections (OPI), Plan-of-Thought (PoT) Backdoor, and Memory Poisoning Attacks, which target the user query, observations, system prompts, and memory retrieval respectively of the agent during action planning and execution.
@@ 2 Related Work
Backdoor Attacks in LLM and LLM Agent. Backdoor attacks embed triggers into the LLMs to generate noxious outputs (Cai et al., 2022; Wan et al., 2023; Li et al., 2024; Zhang et al., 2024a; Wang et al., 2024b). BadChain (Xiang et al., 2024b) has engineered specific trigger words designed to disrupt the Chain-of-Thought (CoT) (Wei et al., 2022) reasoning of LLMs.  Kandpal et al. (2023) utilizes trigger words to disrupt the contextual learning process. Researchers have recently targeted LLM agents for backdoor attacks (Wang et al., 2024a; Yang et al., 2024c; Dong et al., 2024; Hubinger et al., 2024).Wang et al. (2024a); Yang et al. (2024c) contaminates task data for fine-tuning LLM agents, enabling attackers to introduce a threat model. In contrast, the PoT backdoor attack proposed in the paper is a training-free backdoor attack on the LLM agent.
@@ 3.1 Defining Basic Concepts
LLM Agent with Knowledge Bases. We consider LLM agents utilizing knowledge bases, such as RAG for corpus retrieval. For a user query $q$ and its tool list, the agent retrieves relevant memory from a database $D=\{\left(k_{1},v_{1}\right),\ldots,\left(k_{|\mathcal{D}|},v_{|\mathcal{D}|}\right)\}$ of query-solution pairs (Wang et al., 2024c). LLM agents use an encoder $E_{q}$ to map both the query and keys into a shared embedding space. A subset $\mathcal{E}_{K}(q\oplus\mathcal{T},\mathcal{D})\subset\mathcal{D}$ is retrieved, containing the $K$ most relevant keys and values based on the similarity between $q\oplus\mathcal{T}$ and the database keys. Formally, an agent using RAG aims to maximize:
@@ 3.2 Threat Model
Adversarial Goal. Generally, the attacker aims to mislead the LLM agent into using a specified tool, compromising its decision-making in Direct Prompt Injections (DPI), Observation Prompt Injections (OPI), Memory Poisoning, and Mixed Attacks. The Adversarial goal is to maximize:
@@
where the adversary aims to maximize the expected probability that the agent when influenced by adversarial modifications $\theta_{\text{malicious}}$, performs a malicious action $a_{m}$ for a given input query $q$. Apart from this, a Plan-of-Thought (PoT) backdoor attack should keep benign actions for clean queries. Other notations are the same as those in Eq. 1. The Adversarial goal is to maximize:
@@
Adversary’s Background Knowledge and Capabilities. ① Tools. The attacker knows every detail of the attack tools, such as their name and functionality. Moreover, the attacker can integrate their attack tools into the agent’s toolkit, such as manipulating third-party API platforms to add malicious tools, like the RapidAPI platform (Gino, 2024). ② Backbone LLM. The attacker lacks knowledge about the agent’s backbone LLM, including architecture, training data, and model parameters. The agent interacts with the LLM solely through API access, without the ability to manipulate the LLM’s internal components. ③ System Prompts. The attacker can also craft and insert prompts into the agent’s system prompt $p_{\text{sys}}$ to deploy the prompt as a new agent, like through ChatGPT plugins (OpenAI, 2024a). ④ User Prompts. We adopt the common assumption from prior backdoor attacks on LLMs (Kandpal et al., 2023; Cai et al., 2022), which posits that the attacker has access to the user’s prompt and can manipulate it, such as by embedding a trigger. This assumption is realistic in scenarios where users rely on third-party prompt engineering services, which could be malicious, or when a man-in-the-middle attacker (Conti et al., 2016) intercepts the user’s prompt by compromising the chatbot or the input formatting tools. ⑤ Knowledge Database. Unlike previous scenarios with white-box access to RAG databases (Zhong et al., 2023) and RAG embedders (Chen et al., 2024), the attacker has black-box access to RAG databases and embedders.
@@ 4 Formalizing Attacks and Defenses in LLM Agents
As shown in Fig. 1, the LLM agent handles tasks involving system prompts, user prompts, memory retrieval, and tool usage, all of which are vulnerable to attacks. An intuitive method is direct prompt manipulation during the user prompt step, where attackers design malicious prompts to directly call the attack tools (Sec. 4.1.1 DPI Attacks). Tool usage is also at risk due to reliance on third-party platforms that may contain malicious instructions (Sec. 4.1.2 OPI Attacks). Additionally, the memory module can be compromised (Sec. 4.2 Memory Poisoning Attacks), and the hidden system prompt is another attack target, where we propose a PoT-based backdoor attack (Sec. 4.3). These attacks can also be combined into mixed attacks (Sec. 4.3 Mixed Attacks). After that, we define the defenses to the attacks above in Sec. 4.5. Finally, we provide attacking examples in Sec. A.2.2.
@@ 4.1 Formalizing Prompt Injection Attacks
Next, we introduce prompt injection attacks, including DPI, which directly manipulates the agent via user prompts, and OPI, which embeds malicious instructions in tool responses.
@@ 4.1.1 Direct Prompt Injection Attacks
4.1.2 Observation Prompt Injection Attacks
Detailed Adversarial Goal. We define the OPI (Observation Prompt Injection) attack as follows:
Definition 2 - Observation Prompt Injection Attack : Considering an LLM agent provided with a target instruction prompt $q^{t}$, a tool list of all available tools $\mathcal{T}$, a target tool list $\mathcal{T}^{t}\subset\mathcal{T}$ for a target task $t$, it obtains an observation set $\mathcal{O}=(o_{1},\cdots,o_{m})$ from the agent’s task execution trajectory. An OPI attack injects an injected instruction $x^{e}$ of an injected task $e$ to any step $i$ of $\mathcal{O}$, denoted as $\mathcal{O}\oplus x^{e}=(o_{1},\cdots,o_{i}\oplus x^{e},\cdots,o_{m})$, and injects an attack tool list $\mathcal{T}^{e}$ to $\mathcal{T}$, such that the agent performs the injected task apart from the intended target task.
@@ 4.1.3 Attack Framework for Different Prompt Injection Ways
Based on Definitions in Sec. 4.1.1 and Sec. 4.1.2, an adversary injects harmful content into the data $x^{t}$, leading the LLM agent to execute an unintended task $x^{e}$ using an attacker-specific tool. For a DPI attack, $x^{t}$ is the target instruction prompt $q^{t}$. For an OPI attack, $x^{t}$ is an observation result $o_{i}\in\mathcal{O}$, such as a response of an API tool called by the agent in the task execution process. We refer to the data containing this malicious content as compromised data, denoted by $\tilde{x}$. Various prompt injection attacks employ different methods to generate the compromised data $\tilde{x}$, using the original target data $x^{t}$, injected instruction $x^{e}$ of the malicious task. For simplicity, we represent a prompt injection attack with $\mathcal{P}$. Formally, the process to generate $\tilde{x}$ can be described as follows:

### v2

@@ Abstract
Although LLM-based agents, powered by Large Language Models (LLMs), can use external tools and memory mechanisms to solve complex real-world tasks, they may also introduce critical security vulnerabilities. However, the existing literature does not comprehensively evaluate attacks and defenses against LLM-based agents. To address this, we introduce Agent Security Bench (ASB), a comprehensive framework designed to formalize, benchmark, and evaluate the attacks and defenses of LLM-based agents, including 10 scenarios (e.g., e-commerce, autonomous driving, finance), 10 agents targeting the scenarios, over 400 tools, 27 different types of attack/defense methods, and 7 evaluation metrics. Based on ASB, we benchmark 10 prompt injection attacks, a memory poisoning attack, a novel Plan-of-Thought backdoor attack, 4 mixed attacks, and 11 corresponding defenses across 13 LLM backbones. Our benchmark results reveal critical vulnerabilities in different stages of agent operation, including system prompt, user prompt handling, tool usage, and memory retrieval, with the highest average attack success rate of 84.30%, but limited effectiveness shown in current defenses, unveiling important works to be done in terms of agent security for the community. We also introduce a new metric to evaluate the agents’ capability to balance utility and security. Our code can be found at https://github.com/agiresearch/ASB.
@@ 1 Introduction
Large Language Models (LLMs) have rapidly advanced in their capabilities, enabling them to perform tasks such as content generation, question answering, tool calling, coding and many others (Kojima et al., 2024; Jin et al., 2025a; Huang et al., 2022; Jin et al., 2025b; Xu et al., 2024). This has paved the way for developing AI agents that combine LLMs with tools and memory mechanisms capable of interacting with broader environments (Ge et al., 2023a; Xu et al., 2025a). These LLM-based agents have the potential to be deployed in various roles, such as safety-critical domains like financial services (Yu et al., 2023b), medical care (Abbasian et al., 2024; Yang et al., 2024b), e-commerce (Zhu et al., 2024; Xu et al., 2025b) and autonomous driving (Mao et al., 2024). As shown in Fig. 1, an LLM-based agent based on ReAct framework (Yao et al., 2022) usually operates through several key steps when solving a task: ① Defining roles and behaviors via a system prompt. ② Receiving user instructions and task details. ③ Retrieving relevant information from a memory database. ④ Planning based on the retrieved information and prior context. ⑤ Executing actions using external tools.
@@
Although recent research on LLM agents and advanced frameworks has made significant progress, the primary emphasis has been on their effectiveness and generalization (Qin et al., 2024; Mei et al., 2024; Ge et al., 2023b), with their trustworthiness remaining largely under-investigated (Hua et al., 2024). Specifically, while each of these steps mentioned above enables the agent to perform highly complex tasks, they also provide attackers with multiple points of access to compromise the agent system. Each stage is vulnerable to different types of adversarial attacks. Although several benchmarks have been proposed to evaluate the security of LLM agents, such as InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024), they are often limited by their scope, assessing either a single type of attack, i.e., Indirect Prompt Injection, or operating in only a few scenarios, such as financial harm and data security. We compare them with ASB in Sec. B.3 in detail. To address these limitations, we introduce Agent Security Bench (ASB), a comprehensive benchmark that formalizes and evaluates a wide range of adversarial attacks and defenses on LLM-based agents in ten different scenarios.
@@
Primarily, ASB covers various attack and defense types targeting each operational step of LLM-based agents, including system prompt, user prompt handling, tool usage, and memory retrieval. It evaluates Direct Prompt Injections (DPI), Indirect Prompt Injections (IPI), Memory Poisoning, Plan-of-Thought (PoT) Backdoor Attacks, Mixed Attacks, and their defenses, offering the first holistic assessment of LLM agents’ security. In detail, a straightforward way to compromise an agent is through DPI, where attackers directly manipulate the user prompt to guide the agent toward malicious actions. Additionally, the agent’s reliance on external tools introduces further risks, particularly as attackers can embed harmful instructions into tool responses, referred to as IPI. Moreover, the planning phase of LLM agents faces security risks, as long-term memory modules like RAG databases (Lewis et al., 2020) can be compromised by memory poisoning attacks, where adversaries inject malicious task plans or instructions to mislead the agent in future tasks. In addition, since the system prompt is typically hidden from the user, it becomes a tempting target for Plan-of-Thought (PoT) Backdoor Attacks, where attackers embed hidden instructions into the system prompt to trigger unintended actions under specific conditions. Finally, attackers can also combine them to create mixed attacks that target multiple vulnerabilities across different stages of the agent’s operation.
@@
Our key contributions are summarized as follows: ① We design and develop Agent Security Bench (ASB), the first comprehensive benchmark including 10 scenarios (e.g., e-commerce, autonomous driving, finance), 10 agents targeting the scenarios, over 400 tools and tasks for evaluating the security of LLM-based agents against numerous attacks and defense strategies. ② We propose a novel PoT Backdoor Attack, which embeds hidden instructions into the system prompt, exploiting the agent’s planning process to achieve high attack success rates. ③ We formalize and categorize various adversarial threats targeting key components of LLM agents, including DPI, IPI, Memory Poisoning Attacks, PoT Backdoor Attacks, and Mixed Attacks, covering vulnerabilities in system prompt definition, user prompt handling, memory retrieval, and tool usage. ④ We benchmark 27 different types of attacks and defenses on ASB across 13 LLM backbones with 7 metrics, demonstrating that LLM-based agents are vulnerable to the attacks, with the highest average attack success rates exceeding 84.30%. In contrast, existing defenses are often ineffective. Our work highlights the need for stronger defenses to protect LLM agents from sophisticated adversarial techniques. ⑤ We introduce the Net Resilient Performance (NRP) metric to assess the agents’ balance between utility and security and emphasize the importance of performance testing on ASB for selecting suitable backbones for agent applications.
Figure 1: Overview of the LLM Agent Attacking Framework, including Direct Prompt Injections (DPI), Indirect Prompt Injections (IPI), Plan-of-Thought (PoT) Backdoor, and Memory Poisoning Attacks, which target the user query, observations, system prompts, and memory retrieval respectively of the agent during action planning and execution.
@@ 2 Related Work
Backdoor Attacks in LLM and LLM Agent. Backdoor attacks embed triggers into the LLMs to generate noxious outputs (Cai et al., 2022; Wan et al., 2023; Li et al., 2024; Zhang et al., 2024a). BadChain (Xiang et al., 2024b) has engineered specific trigger words designed to disrupt the Chain-of-Thought (CoT) (Wei et al., 2022) reasoning of LLMs.  Kandpal et al. (2023) utilizes trigger words to disrupt the contextual learning process. Researchers have recently targeted LLM agents for backdoor attacks (Wang et al., 2024a; Yang et al., 2024c; Dong et al., 2024; Hubinger et al., 2024).Wang et al. (2024a); Yang et al. (2024c) contaminates task data for fine-tuning LLM agents, enabling attackers to introduce a threat model. In contrast, the PoT backdoor attack proposed in the paper is a training-free backdoor attack on the LLM agent.
@@ 3.1 Defining Basic Concepts
LLM Agent with Knowledge Bases. We consider LLM agents utilizing knowledge bases, such as RAG for corpus retrieval. For a user query $q$ and its tool list, the agent retrieves relevant memory from a database $D=\{\left(k_{1},v_{1}\right),\ldots,\left(k_{|\mathcal{D}|},v_{|\mathcal{D}|}\right)\}$ of query-solution pairs (Wang et al., 2024b). LLM agents use an encoder $E_{q}$ to map both the query and keys into a shared embedding space. A subset $\mathcal{E}_{K}(q\oplus\mathcal{T},\mathcal{D})\subset\mathcal{D}$ is retrieved, containing the $K$ most relevant keys and values based on the similarity between $q\oplus\mathcal{T}$ and the database keys. Formally, an agent using RAG aims to maximize:
@@ 3.2 Threat Model
Adversarial Goal. Generally, the attacker aims to mislead the LLM agent into using a specified tool, compromising its decision-making in Direct Prompt Injections (DPI), Indirect Prompt Injections (IPI), Memory Poisoning, Plan-of-Thought (PoT) backdoor attacks and Mixed Attacks. The Adversarial goal is to maximize:
@@
where the adversary aims to maximize the expected probability that the agent when influenced by adversarial modifications $\theta_{\text{malicious}}$, performs a malicious action $a_{m}$ for a given input query $q$. Apart from this, a PoT backdoor attack should keep benign actions for clean queries. Other notations are the same as those in Eq. 1. The Adversarial goal is to maximize:
@@
Adversary’s Background Knowledge and Capabilities. ① Tools. The attacker knows every detail of the attack tools, such as their name and functionality. Moreover, the attacker can integrate their attack tools into the agent’s toolkit, such as manipulating third-party API platforms to add malicious tools, like the RapidAPI platform (Gino, 2024). ② Backbone LLM. The attacker lacks knowledge about the agent’s backbone LLM, including architecture, training data, and model parameters. The agent interacts with the LLM solely through API access, without the ability to manipulate the LLM’s internal components. ③ System Prompts. The attacker can also craft and insert prompts into the agent’s system prompt $p_{\text{sys}}$ to deploy the prompt as a new agent. This aligns closely with real-world scenarios. Attackers can exploit the increasing reliance on third-party tools or services for prompt engineering, because writing effective system prompts often demands expertise, leading users to outsource to external specialists or tools, such as Fiverr (Fiverr, 2024) or ChatGPT plugins (OpenAI, 2024a). These attackers, posing as prompt engineers or providing prompt optimization tools, can embed backdoors into the system prompts. User Prompts. We adopt the common assumption from prior backdoor attacks on LLMs (Kandpal et al., 2023; Cai et al., 2022), which posits that the attacker has access to the user’s prompt and can manipulate it, such as by embedding a trigger. This assumption is realistic when users rely on third-party prompt engineering services, which could be malicious, or when a man-in-the-middle attacker (Conti et al., 2016) intercepts the user’s prompt by compromising the chatbot or the input formatting tools. ⑤ Knowledge Database. Unlike previous scenarios with white-box access to RAG databases (Zhong et al., 2023) and RAG embedders (Chen et al., 2024), the attacker has black-box access to RAG databases and embedders.
@@ 4 Formalizing Attacks and Defenses in LLM Agents
As shown in Fig. 1, the LLM agent handles tasks involving system prompts, user prompts, memory retrieval, and tool usage, all vulnerable to attacks. An intuitive method is direct prompt manipulation during the user prompt step, where attackers design malicious prompts to directly call the attack tools (Sec. 4.1.1 DPI Attacks). Tool usage is also at risk due to reliance on third-party platforms that may contain malicious instructions (Sec. 4.1.2 IPI Attacks). Additionally, the memory module can be compromised (Sec. 4.2 Memory Poisoning Attacks), and the hidden system prompt is another attack target, where we propose a PoT-based backdoor attack (Sec. 4.3). These attacks can also be combined into mixed attacks (Sec. A.1 Mixed Attacks). After that, we define the defenses to the attacks above in Sec. 4.4. Finally, we provide attacking examples in Sec. A.3.2.
@@ 4.1 Formalizing Prompt Injection Attacks
Next, we introduce prompt injection attacks, including DPI, which directly manipulates the agent via user prompts, and IPI, which embeds malicious instructions in tool responses.
@@ 4.1.1 Direct Prompt Injection Attacks
4.1.2 Indirect Prompt Injection Attacks
Detailed Adversarial Goal. We define the IPI (Indirect Prompt Injection) attack as follows:
Definition 2 - Indirect Prompt Injection Attack : Considering an LLM agent provided with a target instruction prompt $q^{t}$, a tool list of all available tools $\mathcal{T}$, a target tool list $\mathcal{T}^{t}\subset\mathcal{T}$ for a target task $t$, it obtains an observation set $\mathcal{O}=(o_{1},\cdots,o_{m})$ from the agent’s task execution trajectory. An IPI attack injects an injected instruction $x^{e}$ of an injected task $e$ to any step $i$ of $\mathcal{O}$, denoted as $\mathcal{O}\oplus x^{e}=(o_{1},\cdots,o_{i}\oplus x^{e},\cdots,o_{m})$, and injects an attack tool list $\mathcal{T}^{e}$ to $\mathcal{T}$, such that the agent performs the injected task apart from the intended target task.
@@ 4.1.3 Attack Framework for Different Prompt Injection Ways
Based on Definitions in Sec. 4.1.1 and Sec. 4.1.2, an adversary injects harmful content into the data $x^{t}$, leading the LLM agent to execute an unintended task $x^{e}$ using an attacker-specific tool. For a DPI attack, $x^{t}$ is the target instruction prompt $q^{t}$. For an IPI attack, $x^{t}$ is an observation result $o_{i}\in\mathcal{O}$, such as a response of an API tool called by the agent in the task execution process. We refer to the data containing this malicious content as compromised data, denoted by $\tilde{x}$. Various prompt injection attacks employ different methods to generate the compromised data $\tilde{x}$, using the original target data $x^{t}$, injected instruction $x^{e}$ of the malicious task. We simply represent a prompt injection attack with $\mathcal{P}$. Formally, the process to generate $\tilde{x}$ can be described as follows: